OpenStudy (deadshot):

Solve for x: 3 over 3 x plus 1 over quantity x plus 4 equals 10 over 7x ?

OpenStudy (deadshot):

\[\frac{ 3 }{ 3x }+\frac{ 1 }{ x+4 }=\frac{ 10 }{ 7x }\]

OpenStudy (anonymous):

3/3 = 1

OpenStudy (deadshot):

So then

OpenStudy (deadshot):

\[\frac{ 3 }{ 3x } becomes x\]

OpenStudy (anonymous):

1/x + 1/(x+4) = (10/7)*(x) // multiply through by x 1 + x/(x+4) = 10/7 x/(x+4) = 10/7-1 x/(x+4) = 3/7 x = (3/7)(x+4) x = (3/7)x + (12/7) x-(3/7)x = 12/7 (4/7)x = 12/7 multiply both sides by 7/4 x = (12/7)*(7/4) x = 12/4 = ?

OpenStudy (deadshot):

so, x=3?

OpenStudy (deadshot):

or should the sign be reversed, making x=-3?

OpenStudy (anonymous):

x = 3

OpenStudy (deadshot):

okay, thanks!